2 cm3 of a hydrocarbon was mixed in combustion with excess (20 cm3) of oxygen. 16 cm3 of residual gas was obtained after explosion and cooling. The residual gas was passed over a solution of potassium hydroxide and 6 cm3 of the residual gas was absorbed. What is the molecular formula of the hydrocarbon and its IUPAC name??
1
Expert's answer
2020-05-13T14:21:19-0400

If after passing 16cm3 of residue gas pover potassium hydroxide 6cm of gas llremain and 10cm3 of water removed.

Then, "Moles= mass\/rfm"

For carbon=6/2=3

For hydrogen=10/2=5

=C3H5 -Cyclopropyl